OpenStudy (anonymous):

Charlie was carrying around the steps to solve Square root of 6x plus 4 = 8. On his way through the Math Factory, Charlie dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 6x plus 4 = 8. A. Divide by 2 from both sides B. Square root both sides C. Divide by 6 from both sides D. Subtract 4 from both sides E. Subtract 8 from both sides F. Square both sides

OpenStudy (anonymous):

lordamercy who writes this nonsense "divide by 2 from both sides" sounds like it was translated from hungarian

OpenStudy (anonymous):

\[\sqrt{6x+4}=8\] 1) square both sides \[6x+4=64\] 2) subtract \(4\) from both sides \[6x=60\] 3) divide both sides by \(6\) \[x=10\] those are the three steps is the correct order

OpenStudy (anonymous):

whatever math teacher wrote this problem should be ashamed of themselves